Mr.Jacobs is 4 times as old as his son. Five years ago, he was 9 times as old as his son. Find Mr.Jacob's age.
Answer by Maths68(1474)

Let
Present age of Mr. Jacob = j
Present age of Son = s

Mr.Jacobs is 4 times as old as his son.
j=4s............(1)
Five years ago,
Mr. Jacob's age = j-5
Son's age = s-5
Then,he was 9 times as old as his son.
j-5=9(s-5)
j-5=9s-45
j=9s-45+5
j=9s-40............(2)
Put the value of j from (1) to (2)
4s=9s-40
4s-9s=-40
-5s=-40
-5s/-5=-40/-5
s=8
Put the value of s in (1)
j=4s
j=4(8)
j=32

Present age of Mr. Jacob = j = 32 years old
Present age of Son = s = 8 years old